### v2

# InfNeRF: Towards Infinite Scale NeRF Rendering with Space Complexity

Abstract.
The conventional mesh-based Level of Detail (LoD) technique, exemplified by applications such as Google Earth and many game engines, exhibits the capability to holistically represent a large scene even the Earth, and achieves rendering with a space complexity of . This constrained data requirement not only enhances rendering efficiency but also facilitates dynamic data fetching, thereby enabling a seamless 3D navigation experience for users.
In this work, we extend this proven LoD technique to Neural Radiance Fields (NeRF) by introducing an octree structure to represent the scenes in different scales. This innovative approach provides a mathematically simple and elegant representation with a rendering space complexity of , aligned with the efficiency of mesh-based LoD techniques. We also present a novel training strategy that maintains a complexity of . This strategy allows for parallel training with minimal overhead, ensuring the scalability and efficiency of our proposed method. Our contribution is not only in extending the capabilities of existing techniques but also in establishing a foundation for scalable and efficient large-scale scene representation using NeRF and octree structures. Code and checkpoints are available at: https://jiabinliang.github.io/InfNeRF.io/
figure description
1. Introduction
Recent progress in Neural Radiance Fields (NeRF) has significantly advanced the field of photo-realistic novel-view synthesisĀ (Mildenhall etĀ al., 2021; Barron etĀ al., 2021, 2023). Surpassing traditional 3D representations like meshes, NeRF demonstrates a superior capability in generating accurate novel views, particularly for transparent or reflective geometries. However, despite these advancements, a predominant focus of NeRF research has been confined to single objects or limited-scale scenes. The challenge of effectively representing extensive, large-scale scenes remains largely underexplored and unresolved.
Large-scale scene reconstruction has a wide range of applications, such as land surveys, search and rescue operations, construction planning, and navigation. A prime example of its potential can be seen in platforms like Google Earth, enabling users to virtually traverse the globe, which not only showcases the magnificent beauty of natural landscapes but also facilitates the exploration of intricate historic buildings. Extending NeRF to handle large-scale scenes is intriguing and important because it opens up possibilities for more immersive and realistic representations of our world.
1.1. Block-NeRF and Mega-NeRF
Simply scaling up the architecture of NeRF proves insufficient to address this challenge, primarily due to the escalating space complexity. As the scenes expand, it becomes impractical to store the entire model in the memory of standard devices such as mobile phones or consumer PCs. Consequently, dividing the model and storing it in disk or cloud systems becomes necessary.
Towards this direction, recent large-scale NeRF approaches, such as Block-NeRFĀ (Tancik etĀ al., 2022) and Mega-NeRFĀ (Turki etĀ al., 2022) divide the whole neural field equally in space into smaller, more manageable blocks, referred to as grid partition in this paper. This allows the rendering device to load only the data relevant to the cameraās current location. However, these methods encounter challenges in scenarios requiring a birdās eye view of the entire scene. Such a panoramic view plays a crucial role in human interactive navigation. For example, to adjust the position of the camera, it is much more efficient to first zoom out, identify the target location, and then zoom in again, compared to the laborious process of navigating at a constant zoomed-in level. In such scenarios, the existing methods need to load all blocks simultaneously, which can overwhelm the memory capacity of the device. In addition to the memory issue, these approaches are prone to aliasing artifacts that arise from an insufficient sampling rate for high-frequency signals as illustrated in Fig. 1.
1.2. Proposed InfNeRF
Example of multi-resolution octree
Illustration of LoD in InfNeRF
In this paper, we propose InfNeRF, a novel method that addresses the above challenges. The core concept of InfNeRF is inspired by the classical mesh-based Level of Detail (LoD) techniquesĀ (Clark, 1976; Erikson etĀ al., 2001; Ribelles etĀ al., 2010; Tanner etĀ al., 1998). An overview is illustrated in Fig. 2. Specifically, we construct a LoD tree for the target scene, where each node encapsulates a small NeRF with a roughly equivalent number of parameters. The root node provides a coarse representation of the entire scene, while its successive children progressively describe more detailed sub-regions. During rendering, each sampling point is forwarded by a different NeRF, depending on its location and radius. The density and color then are accumulated to determine the pixel color.
As illustrated in Fig. 3, InfNeRF only requires a minimum subset of the nodes in rendering, chosen based on their proximity to the camera and their intersection within the frustum. This significantly reduces the memory burden and the I/O time for retrieving parameters from disk or cloud storage. The memory complexity is notably efficient at , where denotes the total information captured in the reconstruction. In the real world, with , we can represent the Earth down to a resolution of centimeters. In general, InfNeRF is like the combination of Mega-NeRFĀ (Turki etĀ al., 2022) and PyNeRF. It divides the scene in both spatial and scale dimensions while existing approaches typically only consider one of them.
Another remarkable feature of InfNeRF is that the parent nodes in the LoD tree act as a smoother, low-pass-filtered version of the scene. Therefore, the proposed method naturally removes high-frequency components in distant views, which effectively reduces the aliasing artifacts and enhances the visual quality of the generated images (Fig.Ā 1). In our experiments, InfNeRF achieves superior rendering quality for large-scene reconstruction, with an improvement of over 2.4 dB in PSNR while accessing only 17% of the total parameters.
We also propose a pruning algorithm to prune the octree into an effective and compact representation using sparse points from SfM. This algorithm adapts the tree to each scene intelligently without any human intervention.
In addition, the training of InfNeRF only increases slightly and maintains a time and space complexity of . Moreover, this strategy allows for the training process to be effectively parallelized across machines.
InfNeRF imposes no constraints on the underlying NeRF model and makes no assumptions about the scene, making it an adaptable framework that can seamlessly accommodate and leverage the rapid advancements in NeRF models.
In summary, our contributions are as follows:
ā¢
We propose a novel InfNeRF for large-scale scene reconstruction. By integrating neural fields with the classic LoD, InfNeRF achieves efficient rendering with a significantly reduced memory footprint.
ā¢
Our InfNeRF demonstrates a substantial improvement in rendering quality, with over 2.4 dB increase in PSNR, and exhibits superior anti-aliasing capabilities.
ā¢
We develop an efficient and scalable training strategy tailor-made for InfNeRF, addressing the complexities inherent in large-scale neural field training.
2. Related Work
2.1. Photogrammetry
Among the established reconstruction techniques, mesh-based reconstruction methods called photogrammetry have been used widely for it efficiency and robustness in creating digital twins world. Open-source tools, such as COLMAPĀ (Schonberger and Frahm, 2016) and OpenMVSĀ (Cernea, 2020), have seamlessly integrated this workflow and gained widespread adoption in the academic community. Furthermore, commercial software solutions like ContextCapture and DJI TerraĀ (DJI, 2024) have significantly elevated the robustness and efficiency to meet industrial standards.
2.2. Level of Detail
When the scene is large, it becomes impractical to store it entirely in memory. In this case, the Level of Detail (LoD) or Hierarchical Level of Detail (HLoD)Ā (Clark, 1976; Erikson etĀ al., 2001; Hoppe, 1998, 2023; Ulrich, 2002; Luebke, 2003) provides a framework for efficiently storing, fetching, and rendering the scene.
Take tiled 2D mapĀ (Tiled web map, 2024) used widely in 2D navigation system for an example. The maps in different resolutions are cut into 256256 PNG images referred to as tiles. These tiles are organized into a quadtree structure, typically comprising 22 levels. This hierarchical arrangement allows web browsers to rapidly fetch data and render any location at any zoom level with constant space complexity.
A similar technique can also be applied to 3D mesh. After the reconstruction of the scene into one detailed mesh, this mesh undergoes multiple simplifications Ā (Daniels etĀ al., 2008) to reduce the number of triangles. Concurrently, its texture also undergoes downsampling multiple timesĀ (Williams, 1983). Together this results in a mesh pyramid with varying levels of detail, ranging from fine to coarse. Subsequently, these meshes are partitioned into numerous tiles and organized hierarchically, resembling a tree structure similar to a quadtree in 2D mapping. During rendering, 3D engine such as CesiumJSĀ (CesiumJS, 2024) or OpenSceneGraphĀ (Burns and Osfield, 2004) walks the tree, then progressively fetch tiles based on their proximity and intersection with the frustum. This LoD technique enables users to have a seamless 3D navigation experience in a very large scene while maintaining a low memory footprint.
2.3. NeRF
NeRF (Mildenhall et al., 2021) employs Multi-Layer Perceptrons (MLPs) for NeRF to capture a continuous density and viewpoint-dependent radiance of a scene. Methods like TensoRF (Chen et al., 2022) and MERF (Reiser et al., 2023) decompose the dense grid into low dimension feature tensor, while InstantNGP (Müller et al., 2022) opts to store these vectors in a sparse hash table. These strategies greatly enhance model compactness while maintaining speed.
2.4. Large-Scale Scene and Partition at Space Dimension
While previous NeRF approaches primarily focus on scenes of limited scale, scaling up NeRF to handle large-scale scenes, such as cities or even the entire Earth, would empower a broader range of applications. Mip-NeRF 360Ā (Barron etĀ al., 2022) proposed a scene contraction technique to compress the unbounded scene into a bounded region. However, this approach compromises quality in the compressed areas. Grid-Guided NeRFĀ (Xu etĀ al., 2023) extends its capacity by combining a 2D feature grid and MLP, but its reliance on a 2D plane assumption limits its capability to handling complicated 3D scene. Block-NeRFĀ (Tancik etĀ al., 2022), Mega-NeRFĀ (Turki etĀ al., 2022), SMERFĀ (Duckworth etĀ al., 2023) and ScaNERFĀ (Wu etĀ al., 2023) divide the scene into several sub-regions and reconstruct them separately. In its specialized indoor scenario, SMERF can even achieve impressive real-time rendering. However, all these methods suffer from a common issue: when rendering a birdās eye view, the entire model has to be loaded into VRAM, limiting their scalability. Additionally, they may suffer from severe aliasing artifacts.
2.5. Anti-aliasing and Partition at Scale Dimension
Anti-aliasing has been a fundamental problem in computer graphics and image processing for a long time and has been extensively explored in the rendering community. The common approaches to tackle this problem can be categorized into two ways: super-sampling and pre-filtering. In the context of NeRF, super-sampling is casting multiple rays per pixel as in MobileNeRFĀ (Chen etĀ al., 2023), or sampling multiple points with perturbation, as demonstrated in Zip-NeRFĀ (Barron etĀ al., 2023). This straightforward strategy is useful but also computationally expensive. On the other hand, Mip-NeRF Ā (Barron etĀ al., 2021) introduces the concept of pre-filtering through the integrated position encoding. Other research has explored scene partitioning at different scales to mitigate aliasing. For instance, BungeeNeRFĀ (Xiangli etĀ al., 2022) proposes a multi-scale MLP that progressively adds detail when the camera zooms in. Tri-MipRFĀ (Hu etĀ al., 2023a), PyNeRFĀ (Turki etĀ al., 2023) and Mip-VoGĀ (Hu etĀ al., 2023b) train a multi-resolution feature grid. These methods demonstrate impressive anti-aliasing capabilities and consistently high-quality rendering across the entire scene. However, their lack of space partitioning leads to challenges when scaling up, requiring the whole finest layer to render a close-up frame.
In contrast to the methods mentioned in this subsection, the anti-aliasing problem is addressed with inherently no additional cost in InfNeRF. Our parent node in the octree is a downsampled copy of the child nodes. Rendering these nodes at an appropriate zoom level yields a result free from aliasing artifacts.
2.6. Scene Partition with an Octree
Some researches, like PlenOctreesĀ (Yu etĀ al., 2021) and NGLoDĀ (Takikawa etĀ al., 2021) utilize an octree-based approach to represent the scene. PlenOctrees stores the spherical harmonics coefficients in the leaves and uses the octree to accelerate accessing, while in NGLoD, a signed distance field is stored in the octree to represent a LoD closed volume. These direct storage techniques provide fast access and training but require significant memory, limiting their scalability. In contrast, InfNeRF employs many compact hash tables as the underlying data storage. The shallow octree is primarily used for organizing the hash tables into a LoD structure. This approach results in a much more compact LoD representation and much better scalability.
2.7. Gaussian Splatting
In recent times, Gaussian splattingĀ (Kerbl etĀ al., 2023) has gained significant attention due to its fast rendering speed compared to NeRF. Additionally, the introduction of Mip-SplattingĀ (Yu etĀ al., 2024) effectively reduces aliasing artifacts across different resolutions. Then Hierarchical GaussiansĀ (Kerbl etĀ al., 2024), VastGaussianĀ (Lin etĀ al., 2024), and CityGaussianĀ (Liu etĀ al., 2024), explore its potential in large scene reconstruction. While Gaussian splatting may not achieve the same PSNR as traditional NeRFĀ (Liu etĀ al., 2024) in challenging scenarios such as varying light conditions or camera exposure, it remains a promising technique for large-scale scene reconstruction in the future.
3. Method
We present InfNeRF, a framework for efficient large-scale scene reconstruction using a LoD octree structure (SectionĀ 3.1). This tree structure naturally enables anti-aliasing rendering (SectionĀ 3.2), which significantly enhances the visual quality of rendered images. We also introduce tree pruning to improve model sparsity and compactness, along with a rendering approach tailored for the pruned structure (SectionĀ 3.3). Moreover, we propose an efficient training strategy for InfNeRF in SectionĀ 3.4. Finally, we provide a comprehensive analysis of the computational complexity of InfNeRF in supplementary material, offering insights into the efficiency and scalability of our approach in handling large-scale scene reconstructions.
3.1. Tree Structure
The core of our InfNeRF is a LoD structure, essentially an octree where each node represents a specific cubic space of the scene, known as the nodeās Axis-Aligned Bounding Box (AABB). The octree is illustrated in the middle of Fig. 2. InfNeRF begins with the entire scene encapsulated within a root node, which is then divided into eight smaller cubes, corresponding to the child nodes. This division process continues recursively, partitioning the scene into increasingly finer parts, both in terms of space (size of the area) and scale (level of detail).
Each node in this tree is paired with its own NeRF, which represents the nodeās AABB at a certain scale. Our approach is flexible regarding the choice of NeRF model. It can be a simple MLPĀ (Mildenhall etĀ al., 2021), Instant-NGPĀ (Müller etĀ al., 2022), triplaneĀ (Chan etĀ al., 2022; Hu etĀ al., 2023a), or even a combination of them, as long as it can query the density and the color for a given 3D point and viewing direction :
In this work, we use Instant-NGP for its efficiency in training. Instant-NGP discretizes each cube into a grid of voxels, and the number of voxels along each axis of the cube is defined as grid size. Correspondingly, the Ground Sampling Distance (GSD) of a node can be approximated as:
whose unit is in meters. The GSD, in this context, serves as a measure of resolution or granularity, indicating the level of detail that the node can capture about the physical world - the smaller the GSD, the finer the details. For example, if a node represents a 1 km3 cube with a grid size of 2048, the GSD will be meters. As each node in our octree shares the same grid size, the child nodes of have a GSD of meters, thereby capturing more detailed information. This is also illustrated in Fig.Ā 2, where the GSD of a cube at level 1 is half that of a cube at level 0.
3.2. Anti-Aliasing Rendering
Similar to the original NeRF, the rendering process of InfNeRF accumulates the density and color of sampling points along a ray to determine the RGB of a pixel. However, the key difference is that the sampling points are queried from different nodes of our octree as shown in Fig.Ā 2 rather than a single NeRF . In the following, we show how we find the corresponding node for a given sampling point.
According to the sampling theoryĀ (Williams, 1983), a sample on the ray is not merely an infinitely small point, but rather a Gaussian sphere with a certain volume. For simplicity, we model this as a sphere with radius that corresponds to the pixel size,
as illustrated on the left of Fig. 2. Then, given a sampling sphere with position and radius , we aim to find the node whose granularity (GSD) matches and whose AABB contains . In this way, we ensure InfNeRF samples a proper region in the 3D space rather than just a single point, leading to a desirable anti-aliasing effect.
We observe that when the sampling sphere falls between two discrete levels, small aliasing occurs. Nevertheless when sampling points transition between levels, discrepancies in density or color may arise, engendering an undesirable āpopupā effect when camera moves. To mitigate these phenomenons, both PyNeRF and Tri-MipRF employ interpolation techniques between adjacent levels which will double the computation. However, we pursue a simpler approach by introducing stochastic perturbations to the radius of the sampling sphere.
This strategy use human eyeās natural blending effect to attenuate the abrupt transitions without incurring additional computational cost. Finally, the level to sample a given sphere is determined by:
where denotes the GSD of the root node, i.e., level 0.
As illustrated in Fig.Ā 3(a), when the camera is far away from the scene, all spheres are relatively large, and therefore, their densities and colors are estimated by the root node, resulting in a smooth anti-aliasing effect. Conversely, as shown in Fig.Ā 3(b), zooming in with the camera reduces the size of the spheres, which shifts the estimation to the leaf nodes, yielding a sharp and detailed image. As InfNeRF necessitates only a minimum subset of the nodes for rendering a frame, it requires loading merely number of parameters from the whole model into memory. The upper bound is reached when the camera looks at the distant horizon, as illustrated in Fig.Ā 3(c). In this scenario, multiple-level nodes are engaged simultaneously, from the root node representing the blurring horizon to the leaf node representing foreground details.
3.3. Tree Pruning
While we can build a perfect octree as in SectionĀ 3.1, this may not be efficient, particularly for non-uniformly sampled scenes. To optimize the structure, it is important to identify the essential nodes necessary for rendering all input images, effectively pruning the superfluous parts of the tree.
In this work, we utilize sparse points obtained from the Structure from Motion (SfM) to approximate scene geometry. Each sparse point observed in images is treated as spheres, each with a radius as defined in SectionĀ 3.2. We find the corresponding nodes for all the spheres in an ideal, infinitely deep octree as described in SectionĀ 3.2. We retain only the nodes that correspond to sparse spheres and their respective parent nodes, while the others are pruned away.
Compared to Mega-NeRF or Block-NeRF, which split the scene uniformly based on space occupancy, our approach offers a more intelligent resource allocation strategy. It adaptively creates deeper branches in areas with rich details and shallower branches in coarse regions.
This approach is robust as the subtree pruning is only restricted to cases where there are no sparse points in a relatively large cube. As further detailed below, even in the rare case that a subtree is pruned by mistake, like in low texture area, the affected volume can still be well reconstructed by its parent node, albeit in a coarser manner. This ensures that, while fine details may be omitted, the essential structure is preserved.
3.3.1. Pruned Tree Rendering
The pruned tree structure introduces new challenges in InfNeRF rendering, as the optimal node determined by Eq.Ā 5 may no longer exist after pruning. In such scenarios, we instead process the sampling sphere with the nearest available ancestor of the intended node.
Specifically, given a sampling sphere , the rendering process recursively travels down the tree. If the sphere successfully arrives at the node of the expected level, the density and color of the sphere will be estimated by that node. However, if this process is interrupted by a pruned node, the rendering of the sphere reverts to its parent node. We outline this rendering strategy with the pseudo-code in AlgorithmĀ 1. The additional computational overhead introduced by tree traversal includes times floating-point comparisons and memory reads, which can be considered negligible when compared to the NeRF evaluation.
3.4. Training
3.4.1. Training Loss
To train the proposed InfNeRF, we accumulate the density and color of all spheres sampled on a ray using volume renderingĀ (Mildenhall etĀ al., 2021) and compare the resulting RGB values with the pixels of the observed images using L2 loss . To remove the cloudy effect from the unobserved areas, particularly in the sky, and enhance visual quality in the birdās-eye view, we introduce a transparency loss , which uniformly samples points within the AABB and uses L1 loss to encourage their density to approach 0.
To further improve the result, we also incorporate the distortion loss fromĀ (Barron etĀ al., 2022) and a regularization loss who measures the L2 differences of the density between levels, to promote the density consistency.The total loss of our training is summarized as:
3.4.2. Pyramid Supervision.
The images in our datasets are captured by drones at relatively close proximity, resulting in a lack of multi-resolution information necessary to supervise the higher-level nodes. To mitigate this issue, following PyNeRFĀ (Turki etĀ al., 2023)ās approach, we augment the training set using an image pyramid constructed by repeatedly downsampling the high-resolution images. The lower-resolution image with a larger pixel size, create larger sampling spheres according to Eq.Ā 3. These larger sampling spheres are sampled from higher-level nodes in the octree, effectively supervising their training. To account for the pixel imbalance across different pyramid levels, we perform uniform sampling in pixel domain, where a high-resolution image receives four times the sampling rate compared to its low-resolution counterpart. This strategy facilitates more balanced training across different levels of the image pyramid.
3.4.3. Distributed Training
As the scene grows, the demands on VRAM and training time also increases. Straightforwardly using existing distributed training approaches, such as Distributed Data Parallel (DDP), leads to increased VRAM consumption and substantial communication overhead. To address these limitations, we propose a new distributed training strategy.
Given a specific tree level , we split the octree into two parts: the upper tree which includes all nodes from level 0 to level , and the lower part which is a forest including a maximum of subtrees. During training, the upper tree is shared across all devices, and the parameters are updated concurrently. Meanwhile, individual subtrees are assigned among different devices and updated independently. In other words, each device only maintains the shared upper tree and a subset of the lower subtrees, effectively reducing the VRAM footprint and communication overhead. For instance, letās consider a perfect octree with four levels. When the root node is shared , the shared upper tree contains only a tiny fraction, approximately 0.17%, of all the parameters in the entire tree. During forward pass, if a sampling sphere descends to a subtree that is not owned by the current device, it will be forwarded by its parent node in the shared tree as demonstrated in Fig.Ā 4. Consequently, during training, the only data transfer is the all-reduce operation on the gradient of the shared upper tree, which leads to a more efficient and scalable training process.
Illustration of the distributed training in InfNeRF
Furthermore, the training data can also be divided among devices for further acceleration. Intuitively, if a ray on a particular device does not intersect with the AABB of any subtree on that device, it would be not sensible to sample the corresponding pixel of the ray. Therefore, we find a 2D AABB for each device by projecting the sparse points inside the subtreesā 3D AABBs back onto images. Then, during the training phase, only the pixels that fall within the 2D AABB are sampled and utilized. The experimental result of distribution training is provided in supplementary materials.
4. Experiment
InfNeRF is implemented using the NeRFStudio frameworkĀ (Tancik etĀ al., 2023). Each nodeās NeRF is an Instant-NGP with default parameters such as a grid size of 2048 and 16 levels of hash tables, each with a size of . Following NeRF-wĀ (Martin-Brualla etĀ al., 2021), each image is assigned an appearance embedding vector to encode its lighting condition. More implementation details such as learning rate and experiment settings are provided in supplementary materials.
4.1. Experiment On Large-Scale Dataset
4.1.1. Dataset
To evaluate the scalability of our method, we adopt a similar approach to Mega-NeRF (Turki etĀ al., 2022), by conducting our main experiment using four real-world large-scale urban scenes dataset captured by drone from Mill 19(Turki etĀ al., 2022) and UrbanScene3D (Lin etĀ al., 2022). Each dataset consists of around two thousand 4K images, amounting to 40 gigapixels and spanning 1 km. Notably, they are approximately 200 times larger than the scenes reported in other general NeRF work. Our octree is limited to 4 levels, resulting a reasonable GSD ranges from 0.4 m of the root node to 5 cm of the leaf node. The total model size is about 3 GB.
4.1.2. Result on Space Complexity
The paramount objective of our study is to reduce the required model size in rendering, which refers to the sum of all sub NeRFsā size required for rendering one frame. It is the total information that needs to be retrieved from the clouds and stored in the local device during real-time rendering. To demonstrate the impact of our octree on the required model size, we compare InfNeRF with three other methods with similar total model sizes but with different architectures. These three methods are: (1) NerfactoĀ (Tancik etĀ al., 2023), (2) Nerfacto-s, a 4 levels Nerfacto, representing the scale partition methods like PyNeRF (Turki etĀ al., 2023), (3) InfNeRF leaf, an InfNeRF without upper tree structure, representing the space partition methods like Mega-NeRFĀ (Turki etĀ al., 2022).
In this experiment, we examine the rendering of a zoom-out trajectory, starting from a close-up view and transitioning to a birdās-eye view in VGA resolution. The trajectory and the corresponding rendered images are shown in Fig. 1. We present the required model size in megabytes for six keyframes in Fig.Ā 5. When rendering the close-up view, the space partition method InfNeRF leaf proves advantageous as it only requires a single node of the scene and has low space requirements. In contrast, the scale partition methods Nerfacto-s necessitate the detailed version of the entire scene, resulting in a higher space complexity. As the camera moves away, the space partition method InfNeRF leaf now requires all blocks to render the complete scene, while the scale partition method reduces its requirement by only rendering the low-resolution node. However, InfNeRF consistently exhibits its low space complexity in both scenarios, with a memory footprint of only 17% compared to InfNeRF leaf and 12.3% compared to Nerfacto-s.
4.1.3. Result on Rendering Quality
As zoom-out is a crucial use case in our paper, we not only report metrics at the finest resolution, denoted as PSNR0, but more importantly, the mean metrics over 6 resolution levels, denoted as PSNR, SSIM, LPIPS. We compare our approach with (1) NerfactoĀ (Tancik etĀ al., 2023), whose model size is scaled up to InfNeRF, (2) PyNeRFĀ (Turki etĀ al., 2023), also being scaled up to InfNeRF, (3) Mega-NeRFĀ (Turki etĀ al., 2022) with 25 partitions, (4) InfNeRF compact (InfNeRF-c), whose model size is similar to Mega-NeRF.
Scene | Residential | Sci-Art | Rubble | Building | ||||||||||||
Metric | PSNR0 | PSNR | SSIM | LPIPS | PSNR0 | PSNR | SSIM | LPIPS | PSNR0 | PSNR | SSIM | LPIPS | PSNR0 | PSNR | SSIM | LPIPS |
Nerfacto | 24.32 | 24.31 | 0.772 | 0.216 | 26.75 | 26.48 | 0.827 | 0.182 | 24.99 | 26.09 | 0.692 | 0.336 | 22.53 | 22.94 | 0.663 | 0.325 |
PyNeRF | 21.30 | 24.13 | 0.753 | 0.299 | 23.85 | 24.85 | 0.815 | 0.242 | 23.48 | 27.08 | 0.744 | 0.347 | 21.70 | 24.93 | 0.743 | 0.279 |
Mega-NeRF | 22.52 | 22.85 | 0.748 | 0.265 | 24.37 | 24.22 | 0.808 | 0.202 | 24.17 | 26.02 | 0.73 | 0.294 | 21.70 | 22.43 | 0.696 | 0.282 |
InfNeRF-c | 23.84 | 24.87 | 0.785 | 0.198 | 25.83 | 25.81 | 0.805 | 0.212 | 24.62 | 27.27 | 0.731 | 0.274 | 22.52 | 24.02 | 0.703 | 0.280 |
InfNeRF | 24.66 | 25.46 | 0.820 | 0.158 | 27.03 | 26.96 | 0.839 | 0.192 | 25.12 | 27.92 | 0.781 | 0.203 | 23.11 | 24.70 | 0.740 | 0.242 |
The PSNR0 (full resolution) and all 3 metrics for multi-resolution are listed in Table 1. The mean across the four datasets are also plotted in Fig. 6. When comparing two models with similar model sizes, InfNeRF and Nerfacto, theoretically, they should exhibit similar PSNR0. Our results confirm this hypothesis. As depicted along the horizontal axis of Fig. 6, InfNeRF slightly outperforms Nerfacto by 0.3 dB, demonstrating that InfNeRFās octree partitioning and the limited memory requirement do not compromise the rendering quality. Furthermore, when considering results of multi-resolution rendering, ploted along the vertical axis, InfNeRF outperforms Nerfacto by 1.31 dB, PyNeRF by 1.01 dB and even outperform Mega-NeRF by 2.4 dB. This illustrates the crucial role of the LoD octree in anti-aliasing rendering. The same conclusion can be drawn from the quality comparison in Fig. 9. In the last three rows of each dataset, which display the low-resolution images, we can observe that without the anti-aliasing effect, Nerfacto and Mega-NeRF generate noisy results, particularly evident in the heavy moire pattern in the left bottom dataset solar panel region. In contrast, InfNeRF generates satisfactory results at both fine and coarse levels, showcasing its superiority in handling multi-resolution rendering. More detailed images are provided in Fig.Ā 10.
4.2. Experiment On MipNeRF 360 Dataset
In addition to our previous large-scale experiment, we conduct another experiment with the garden scene of MipNeRF 360 Dataset. This smaller-scale scene allows us to evaluate and compare the performance of various methods. The following methods were included in our comparison: (1) Gaussian Splatting(Kerbl etĀ al., 2023). (2) Mip-Splatting(Yu etĀ al., 2024). (3) NerfactoĀ (Tancik etĀ al., 2023). (4) PyNeRFĀ (Turki etĀ al., 2023). All NeRF based methods have similar model size. More details of the experiment setting can be found in the supplementary material.
Resolution | |||||
Gaussian Splatting | 26.52 | 25.18 | 21.58 | 19.01 | 17.266 |
Mip-Splatting | 26.87 | 27.85 | 28.83 | 28.69 | 27.76 |
Nerfacto | 23.18 | 23.5 | 24.08 | 23.32 | 22.09 |
PyNeRF | 23.79 | 24.76 | 27.04 | 27.63 | 24.82 |
InfNeRF | 26.27 | 27.04 | 28.2 | 28.87 | 27.23 |
(1) w/o Trans Loss | 25.63 | 26.45 | 27.54 | 27.92 | 26.18 |
(2) w/o Perturbation | 25.90 | 26.87 | 27.86 | 28.23 | 27.20 |
(3) w/o Upper Tree | 24.71 | 25.22 | 24.95 | 24.02 | 22.69 |
(4) w/o Pyramid | 26.12 | 23.04 | 13.00 | 8.98 | 8.52 |
(5) w 3 levels | 24.78 | 25.2 | 26.2 | 27.26 | 25.96 |
As shown in Table. 2, InfNeRF exhibits competitive outcomes in high-resolution when compared to 3DGSĀ (Kerbl etĀ al., 2023) and Mip-SplattingĀ (Yu etĀ al., 2024). It surpasses PyNeRF and Nerfacto by an margin of 3 dB. As the resolution decreases, 3DGSĀ (Kerbl etĀ al., 2023) and Nerfactoās performance start to drop. Although anti-aliasing may not be the primary focus of InfNeRF, it consistently delivers commendable results without any additional computational cost.
4.3. Ablation Study
Table. 2 also includes an ablation study of our model on garden scene. (1) The transparency loss, improves slightly PSNR, and more importantly, removes the fog blocking the bird eyes view, as shown in Fig. 8. (2) Sampling radius perturbation not only smooth the transition between different LOD levels in some dataset, as depicted in Fig. 7, but also improve the overall PSNR. (3) Training without the upper tree structure, or (4) training without image pyramid, results in a significant reduction in PSNR at the low-resolution. (5) Removing one level of the tree reduces the model size by 4, and decreases the performance by 2 dB. (6) Tree pruning reduces the size of the tree to only 19 nodes, which accounted for only 26% of the full octree.
5. Conclusion and Future work
In this work, we introduce InfNeRF for large-scale scene reconstruction. InfNeRF employs an LoD tree to divide the scene in space and scale into cubes, which are reconstructed using many NeRFs. During rendering, InfNeRF selectively retrieves a minimal subset of nodes, significantly reducing memory requirements and I/O time for parameter retrieval from disk or cloud storage and also reducing aliasing artifacts. In our experiments, InfNeRF achieves superior rendering quality, with an improvement of over 2.4 dB in PSNR while accessing only 17% of the total parameters.
InfNeRF demonstrates its potential for large-scale scene reconstruction. However, there are several aspects for future exploration and improvement. Firstly, the reconstruction time and computational burden still exceed traditional, well-optimized photogrammetry methodsĀ (Schonberger and Frahm, 2016; DJI, 2024), necessitating further performance enhancements. Secondly, exploring the combination of octree from diverse image sources, such as satellites, planes, and drones, at different times and scales, could enable the creation of larger scenes. With these works, digitizing the Earth using NeRF will become a compelling possibility for the future.